OpenStudy (anonymous):

A multiple choice has 10 questions. Each question has four answer choices. what is the probability a student randomly guesses the answers and gets exactly six questions correct?

OpenStudy (anonymous):

i think i know how to sove this problem. but im not positive if im correct

ganeshie8 (ganeshie8):

he can get exactly 6 questions correct in \(^{10}C_6 \)ways

OpenStudy (anonymous):

how would i show work?

ganeshie8 (ganeshie8):

and each way has a probability of : \(\large (\frac{1}{4})^6(\frac{3}{4})^4\)

OpenStudy (anonymous):

then i got .25^6 * .75^4

OpenStudy (anonymous):

and then when i began to simplify further. the decimals didnt seem to make sense to me

ganeshie8 (ganeshie8):

so, the total probability for getting "Exactly 6 questions correct" would be : \(^{10}C_6 \large (\frac{1}{4})^6(\frac{3}{4})^4 \)

OpenStudy (anonymous):

isnt that just the equation?

OpenStudy (anonymous):

so the probability would be 1.62%?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

probability is always between [0, 1]

OpenStudy (anonymous):

so it wouldnt be 1.62?

ganeshie8 (ganeshie8):

oh you're saying 1.62 % , then fine :)

OpenStudy (anonymous):

so would that be correct?

ganeshie8 (ganeshie8):

you're right ! probability is 1.62%

ganeshie8 (ganeshie8):

which is same as 0.016

OpenStudy (anonymous):

thank you! theres another part of this problem that i need help with

ganeshie8 (ganeshie8):

shoot

OpenStudy (anonymous):

is getting exactly 10 questions correct the same probability as getting exactly zero correct

OpenStudy (anonymous):

1/4 is the probability of getting a question correct 3/4 is the probability of getting a question incorrect so would i use the same equation from above but replace the 6 with a 10 as well as a 0?

ganeshie8 (ganeshie8):

yes u can, but this is realitively easier than that

ganeshie8 (ganeshie8):

since randomly picking answers are independent events, u can simply multiply the probabilities

ganeshie8 (ganeshie8):

P("getting all 10 questions correct") = \(\large (\frac{1}{4})^{10}\)

ganeshie8 (ganeshie8):

P("getting 0 questions correct") = \(\large (\frac{3}{4})^{10}\)

ganeshie8 (ganeshie8):

are they same ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i dont believe so

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

correct, so they're not same

OpenStudy (anonymous):

okay, and one more part if you dont mind describe the steps needed to calculate the probability of getting at least six questions correct if the student randomly guesses

ganeshie8 (ganeshie8):

atleast 6 questions correct means, he can get 6 or more question correct, right ?

OpenStudy (anonymous):

correct

ganeshie8 (ganeshie8):

So, P("atleast 6 questions correct") = P("Exactly 6 questions correct") + P("Exactly 7 questions correct") + P("Exactly 8 questions correct") + P("Exactly 9 questions correct") + P("Exactly 10 questions correct")

ganeshie8 (ganeshie8):

Find the probability for each event and add them ^^

ganeshie8 (ganeshie8):

you dont need to find it literally, just explain in words that... u need to find all probabilities and add them, okay ?

OpenStudy (anonymous):

okay thank you !

ganeshie8 (ganeshie8):

np.. u wlc :)